# Improving Zero-noise Extrapolation for Quantum-gate Error Mitigation using a Noise-aware Folding Method

### v1

@@ V Discussion
We applied other extrapolation methods, namely, exponential fit [22], adaptive-exponential fit from [30], and Richardson’s method [10], to ZNE execution on the above-mentioned circuit folding methods. However, our benchmarking results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting. Indeed, the models were often degraded by extreme overfitting or underfitting, leading to nonconvergence in some instances. Although improvements from those of linear fit extrapolation were sporadically observed, the results of these alternative fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms.
Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed approach. Therefore, the efficacy and discriminative power of these extrapolations in distinguishing the performances of different extrapolation techniques is questionable. To evaluate our noise-aware folding method against the existing approaches, we conducted comprehensive experiments using linear fit extrapolation. Simulations on various full-noise models demonstrated that our proposed method consistently and efficiently performs across different quantum circuit complexities. Notably, it maintained high extrapolation accuracy as the circuit complexity increased.
Real quantum computer executions also delivered promising results, particularly on smaller circuits. However, at higher qubit counts, the expectation values fell more rapidly with circuit scaling than in the simulation results, affecting the reliability of the extrapolated results. The disparity between the simulation and real quantum computer executions highlights the challenges posed by the noise models, which do not capture all noise sources in actual hardware.
VI Conclusion
@@
Our noise-aware folding method more effectively mitigates quantum errors during computations than the existing methods. By dynamically adjusting the error rates based on hardware noise models, it promises to improve the precision and reliability of quantum computations. The fidelity improvements over the compared methods reached 35% and 26% on quantum computer simulators and real quantum computers, respectively. Although demonstrating robustness in simulations and smaller-scale executions on real quantum computers, the proposed method was challenged by scaling to larger circuits, necessitating further exploration and optimization to broaden its applicability. Owing to its adaptability and versatility, the proposed method can potentially bridge the gap between theoretical expectations and practical outcomes in QC. Further refinements and adaptations are imperative to enhance scalability and reliability, propelling quantum error mitigation methods towards realizing error-free quantum computations.
Acknowledgements.
@@
This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2020-0-00014, A Technology Development of Quantum OS for Fault-tolerant Logical Qubit Computing Environment).

### v2

@@ 5 Discussion
We applied other extrapolation methods, namely, exponential fit [22], adaptive-exponential fit from [30], and Richardson’s method [10], to ZNE execution on the above-mentioned circuit folding methods. However, the results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting model. In fact, the models were often degraded by extreme over-fitting or under-fitting, leading to non-convergence in some instances. Although improvements from those of linear fit extrapolation were sometimes observed, the results of these fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms where a consistent result is required for many iterations. Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed noise-aware folding approach. Hence, the effectiveness of these extrapolations in discerning the performances of various extrapolation techniques are uncertain. Although linear fit extrapolation, like any other models, is unable to extrapolate to zero noise as theoretically proposed due to strong noise in the quantum computer, the consistent improvement in results we observe using linear fit sets it apart from other extrapolation models. However, utilizing zero-noise extrapolation comes with a trade-off of executing more quantum circuits, typically ranging from $2x$ to $5x$ the original amount, depending on the number of $\lambda$ values desired by the user.
Moreover, while QEM techniques like ZNE are primarily employed to enhance the results of NISQ devices, they may also prove practical for use in quantum error correction codes to protect qubits from errors. By employing both error correction codes and mitigation strategies, a more resource-efficient method can be achieved compared to traditional error correction codes alone. In this hybrid approach, QEM techniques can be used to pre-process the noisy qubits before encoding them into a QEC code. This can reduce the error rates of the qubits, making the subsequent QEC code more effective and efficient. For instance, ZNE can be applied to suppress errors in the qubits before encoding them into a surface code or a concatenated code. Alternatively, QEM techniques can be integrated directly into the QEC code, allowing for real-time error mitigation during the execution of the quantum algorithm. This can be particularly useful for codes that are prone to error accumulation, such as those used in quantum simulations or machine learning applications. As advancements in quantum computer technology reduce noise levels, making error mitigation strategies more effective, this potential becomes even more viable. With lower noise levels, QEM techniques can provide a greater reduction in error rates, further enhancing the performance of the hybrid approach.
6 Conclusion
@@
We proposed a noise-aware folding method to improve the zero-noise extrapolation method by employing the noise-adaptive compilation and a folding mechanism that distributes noise uniformly using calibration data based on the quantum hardware noise model. By dynamically adjusting the error rates based on hardware noise models, it improves the precision and reliability of the linear fit extrapolation method results in improving the quantum computer result fidelity The fidelity improvements over the compared methods reached 35% and 31% on quantum computer simulators and real quantum computers, respectively. Although demonstrating robustness in simulations and smaller-scale executions on real quantum computers, the proposed method was challenged by scaling to larger circuits, necessitating further exploration and optimization to broaden its applicability. Further refinements and adaptations are essential to enhance scalability and reliability, driving quantum error mitigation methods towards realizing error-free quantum computations.
[…]
@@
Acknowledgments This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2020-0-00014, A Technology Development of Quantum OS for Fault-tolerant Logical Qubit Computing Environment).
Appendix A ibmq_mumbai’s Properties
This supplementary document provides the experimental setup detail of the quantum computer that we use in our demonstration.
Figure 7: ibmq_mumbai’s qubit connectivity topology.
Qubit T1 (us) T2 (us) Frequency (GHz) Anharmonicity (GHz)
0 111.6961464 179.1588076 5.070549646 -0.328451763
1 122.2484892 197.5837264 4.929863389 -0.331291977
2 81.90697735 155.8248129 4.670281392 -0.336876757
3 167.7648109 122.141877 4.889402396 -0.331179264
4 42.05229013 29.68056521 5.020938057 -0.330174519
5 101.1535787 122.3569753 4.969532374 -0.329757062
6 103.291911 45.33644272 4.965898301 -0.329305699
7 67.91110372 48.58658623 4.894369882 -0.331001624
8 87.45290954 142.5387495 4.79164979 -0.332559138
9 96.66829381 73.24259076 4.954976802 -0.330638393
10 110.9073585 132.0053088 4.958620128 -0.330907906
11 88.46621356 202.3471754 4.666070269 -0.332629168
12 141.6272911 204.5251318 4.74291957 -0.332978886
13 117.1107884 151.0626282 4.88858331 -0.328060893
14 84.32426156 153.5437913 4.780232309 -0.332554813
15 97.45073523 90.18823679 4.8581701 -0.333242963
16 104.390153 262.8877462 4.98035969 -0.329920963
17 104.1706916 85.0072165 5.003339003 -0.329925892
18 122.1904549 212.2866827 4.781291798 -0.333097806
19 117.3925825 251.3554841 4.809727654 -0.33209643
20 130.2027001 214.4300921 5.047734483 -0.328048366
21 109.9161771 200.7424141 4.942677405 -0.331299817
22 130.3648863 211.195153 4.911344811 -0.331838149
23 90.39455816 163.0951773 4.892629733 -0.331514799
24 141.7988532 29.62637317 4.671041833 -0.335931163
25 148.610822 128.6915538 4.758538084 -0.333550575
26 72.49759913 135.6307972 4.954268232 -0.329535731
Table 1: ibmq_mumber’s noise model captured on 26-March-2024 from IBM.
Qubit Readout assignment error Prob meas0 prep1 Prob meas1 prep0 Readout length (ns) ID error Sx error Pauli-X error CNOT error
0 2.26E-02 3.54E-02 9.80E-03 3512.888889 2.74E-04 2.74E-04 2.74E-04 0_1:5.62E-03
1 1.58E-02 2.62E-02 5.40E-03 3512.888889 1.93E-04 1.93E-04 1.93E-04 1_2:8.25E-03;1_4:5.29E-03;1_0:5.62E-03
2 1.15E-02 1.60E-02 7.00E-03 3512.888889 1.16E-04 1.16E-04 1.16E-04 2_3:5.11E-03;2_1:8.25E-03
3 1.15E-02 1.74E-02 5.60E-03 3512.888889 2.33E-04 2.33E-04 2.33E-04 3_2:5.11E-03;3_5:7.57E-03
4 2.73E-02 3.72E-02 1.74E-02 3512.888889 2.77E-04 2.77E-04 2.77E-04 4_7:8.49E-03;4_1:5.29E-03
5 1.74E-02 2.32E-02 1.16E-02 3512.888889 2.33E-04 2.33E-04 2.33E-04 5_3:7.57E-03;5_8:2.25E-02
6 1.72E-02 2.58E-02 8.60E-03 3512.888889 2.57E-04 2.57E-04 2.57E-04 6_7:1.22E-02
7 1.88E-02 3.12E-02 6.40E-03 3512.888889 4.94E-04 4.94E-04 4.94E-04 7_10:8.76E-03;7_6:1.22E-02;7_4:8.49E-03
8 1.97E-02 2.96E-02 9.80E-03 3512.888889 1.90E-04 1.90E-04 1.90E-04 8_11:6.50E-03;8_5:2.25E-02;8_9:1.21E-02
9 1.52E-02 2.34E-02 7.00E-03 3512.888889 2.63E-04 2.63E-04 2.63E-04 9_8:1.21E-02
10 2.28E-02 3.38E-02 1.18E-02 3512.888889 4.45E-04 4.45E-04 4.45E-04 10_12:1.59E-02;10_7:8.76E-03
11 4.44E-02 5.60E-02 3.28E-02 3512.888889 2.47E-04 2.47E-04 2.47E-04 11_14:5.62E-03;11_8:6.50E-03
12 2.83E-02 1.90E-02 3.76E-02 3512.888889 1.62E-04 1.62E-04 1.62E-04 12_13:7.55E-03;12_10:1.59E-02;12_15:4.46E-03
13 1.16E-02 1.46E-02 8.60E-03 3512.888889 1.70E-04 1.70E-04 1.70E-04 13_12:7.55E-03;13_14:6.55E-03
14 4.60E-02 3.74E-02 5.46E-02 3512.888889 1.61E-04 1.61E-04 1.61E-04 14_11:5.62E-03;14_13:6.55E-03;14_16:5.07E-03
15 2.01E-02 2.54E-02 1.48E-02 3512.888889 1.93E-04 1.93E-04 1.93E-04 15_18:5.09E-03;15_12:0.0044594326370586135
16 1.28E-02 2.16E-02 4.00E-03 3512.888889 1.94E-04 1.94E-04 1.94E-04 16_19:0.01074071841628857;16_14:5.07E-03
17 1.98E-02 3.18E-02 7.80E-03 3512.888889 3.48E-04 3.48E-04 3.48E-04 17_18:1.14E-02
18 8.00E-02 9.62E-02 6.38E-02 3512.888889 1.93E-04 1.93E-04 1.93E-04 18_21:6.97E-03;18_15:5.09E-03;18_17:1.14E-02
19 2.31E-01 2.33E-01 2.29E-01 3512.888889 2.49E-04 2.49E-04 2.49E-04 19_22:5.13E-03;19_20:8.25E-03;19_16:1.07E-02
20 1.13E-01 3.26E-02 1.94E-01 3512.888889 1.81E-04 1.81E-04 1.81E-04 20_19:8.25E-03
21 1.91E-02 2.80E-02 1.02E-02 3512.888889 3.70E-04 3.70E-04 3.70E-04 21_18:6.97E-03;21_23:7.72E-03
22 2.33E-02 3.96E-02 7.00E-03 3512.888889 1.87E-04 1.87E-04 1.87E-04 22_19:5.13E-03;22_25:7.16E-03
23 6.02E-02 7.44E-02 4.60E-02 3512.888889 2.62E-04 2.62E-04 2.62E-04 23_21:7.72E-03;23_24:9.50E-03
24 2.17E-02 3.36E-02 9.80E-03 3512.888889 3.61E-04 3.61E-04 3.61E-04 24_23:9.50E-03;24_25:5.22E-03
25 1.31E-02 1.92E-02 7.00E-03 3512.888889 1.42E-04 1.42E-04 1.42E-04 25_22:7.16E-03;25_26:4.25E-03;25_24:5.22E-03
26 1.41E-02 1.90E-02 9.20E-03 3512.888889 1.65E-04 1.65E-04 1.65E-04 26_25:4.25E-03
Table 2: Extended ibmq_mumber’s noise model including full detail of the gate error rate captured on 26-March-2024 from IBM.